# Stepwise Alignment for Constrained Language Model Policy Optimization

### v2

@@ Abstract
Safety and trustworthiness are indispensable requirements for real-world applications of AI systems using large language models (LLMs). This paper formulates human value alignment as an optimization problem of the language model policy to maximize reward under a safety constraint, and then proposes an algorithm, Stepwise Alignment for Constrained Policy Optimization (SACPO). One key idea behind SACPO, supported by theory, is that the optimal policy incorporating reward and safety can be directly obtained from a reward-aligned policy. Building on this key idea, SACPO aligns LLMs step-wise with each metric while leveraging simple yet powerful alignment algorithms such as direct preference optimization (DPO). SACPO offers several advantages, including simplicity, stability, computational efficiency, and flexibility of algorithms and datasets. Under mild assumptions, our theoretical analysis provides the upper bounds on optimality and safety constraint violation. Our experimental results show that SACPO can fine-tune Alpaca-7B better than the state-of-the-art method in terms of both helpfulness and harmlessness.
@@ 1 Introduction
Safe RLHF [15] is a pioneering approach for introducing the (constrained) safe RL paradigm into the alignment of LLMs. As with the standard RLHF pipeline, Safe RLHF trains separate reward and safety models from the human preference datasets and then employs RL to optimize an LM policy. This approach facilitates the acquisition of LLMs that strike a well-balanced compromise between reward (i.e., helpfulness) and safety (i.e., harmlessness). However, the Safe RLHF pipeline is inherently more complex than standard RLHF, as it necessitates 1) fitting separate reward and safety models to preference data and then 2) learning a policy via PPO Lagrangian [36] that simultaneously optimizes an additional parameter (i.e., Lagrangian multiplier) to balance helpfulness and harmlessness. In addition, Safe RLHF often suffers from an issue called exaggerated safety behaviors [9], which results in the model generating harmless but unhelpful responses.
Our contributions. We propose an algorithm called Stepwise Alignment for Constrained Policy Optimization (SACPO) for human value alignment of LLMs while incorporating decoupled reward and safety metrics. As shown in Figure 1, SACPO is a stepwise approach that sequentially aligns an LLM with one metric (e.g., reward) and subsequently with another (e.g., safety). Our theoretical analysis allows us to employ simple RL-free alignment algorithms such as direct preference optimization (DPO, [35]) or Kahneman-Tversky optimization (KTO, [19]) for each alignment without necessitating explicit reward or safety modeling. In a theoretically justified way, SACPO allows us to use different alignment algorithms or parameters for each alignment, thereby enhancing the flexibility of the format or volume of the datasets. To enhance the practicality of SACPO, we further propose an efficient approach called P-SACPO using model merging to balance the trade-off between reward and safety performance. We provide theoretical results on the optimality and safety of the LM policy of SACPO under mild assumptions. Finally, our experimental results show that SACPO can fine-tune Alpaca-7B better than Safe RLHF in terms of both helpfulness and harmlessness.
Figure 1: Safe RLHF [15] respectively fits reward and safety models to reward and safety datasets with human preferences, and then leverages PPO Lagrangian to optimize an LM policy and a Lagrangian multiplier to balance helpfulness and harmlessness. In contrast, SACPO first aligns an LM policy with the reward metric and then realigns the resulting reward-aligned policy with the safety metric (or vice versa). In this process, we can use simple RL-free algorithms (e.g., DPO, KTO) for each step, which leads to simplicity, stability, and flexibility.
@@ 2.3 Safe and Multi-objective Learning from Human Feedback
Safe RLHF. To improve the helpfulness and harmlessness of LLMs, Dai et al. [15] introduce a safety function $g^{\star}$ and then formulate the LLM alignment problem as a policy optimization task of maximizing the reward $r^{\star}$ under a safety constraint. They propose Safe RLHF that first trains reward and safety models (i.e., $r_{\psi}$ and $g_{\psi}$) using two datasets containing reward and safety information, and then solves the following problem using a popular safe RL algorithm called PPO-Lagrangian [36].
@@ 4 Stepwise Alignment for Constrained Policy Optimization
$\max_{\pi}\min_{\lambda\geq 0}\ L(\pi,\lambda,\beta)=\mathbb{E}_{\rho,\pi}\bigl{[}r^{\star}(x,y)+\lambda g^{\star}(x,y)\bigr{]}-\beta\mathbb{D}_{\mathrm{KL}}[\,\pi_{\theta}(y\mid x)\,\|\,\pi_{\mathrm{ref}}(y\mid x)\,]-\lambda b.$ (10)
@@ 4.1 Optimal Policy Can be Directly Obtained from Reward-aligned Policy
For the proof, see Appendix D. Intuitively, Theorem 1 states that we do not have to align a policy for multiple metrics simultaneously and thus we can sequentially align the policy stepwise. Specifically, (11) means that the optimal policy $\pi^{\star}$ is identical to the one obtained by realignment of $\pi^{\star}_{r^{\star}}$ for the safety function $g^{\star}$ with a parameter $\beta/\lambda^{\star}$. Thus, (11) justifies realigning the reward-aligned model $\pi^{\star}_{r^{\star}}$ with respect to the safety function $g^{\star}$. After a simple mathematical transformation of (11) and the fact that $\log Y(x)$ neither depend on $y$ nor $\pi$ for all $x\in\mathcal{X}$, we then have
@@
Advantages of SACPO. By taking a stepwise approach, we can enjoy practical benefits. Let us highlight three major advantages of SACPO. The first advantage is the flexibility of alignment algorithms (e.g., DPO or KTO) and datasets. In practice, depending on the metric, appropriate human feedback should be different (e.g., paired vs. unpaired). SACPO takes a stepwise approach, which allows us to use different algorithms, parameters, or datasets for each metric. Second, we can evaluate the resulting LM policy after each alignment regarding the target metric. This process enables us to prevent starting over the alignment from the beginning. Finally, SACPO justifies us to realign pre-aligned LLMs with our desired metric. This property is practically desirable because we now have easy access to high-quality, open-source LLMs that have been already aligned.
@@ 5 Theoretical Results
The reward and safety functions are parameterized by $\widehat{r}(x,y)=\langle w_{r},\phi(x,y)\rangle$ and $\widehat{g}(x,y)=\langle w_{g},\phi(x,y)\rangle$ for a shared feature mapping function $\phi:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}^{d}$. In addition, the true reward and safety functions satisfy $r^{\star}(x,y)=\langle w_{r}^{\star},\phi(x,y)\rangle$ and $g(x,y)=\langle w_{g}^{\star},\phi(x,y)\rangle$ for some $w_{r}^{\star},w_{g}^{\star}\in\mathbb{R}^{d}$. For regularization, we additionally assume $\|\phi(x,y)\|\leq 1$ for any $(x,y)\in\mathcal{X}\times\mathcal{Y}$ and $\max\{\|w_{r}\|,\|w_{g}\|\}\leq B$.
@@
Theorems 2 and 3 suggest that, regarding both optimality and safety violation, the performance degradation compared to $\pi^{\star}$ is exponential to $\frac{2}{\beta}\,\widehat{\Gamma}_{\mathcal{D}}(x,y,\lambda^{\star})=\frac{2}{\beta}\big{(}\alpha\,\bigl{(}\mathcal{U}_{\mathcal{D}_{r}}(x,y)+\lambda^{\star}\,\mathcal{U}_{\mathcal{D}_{g}}(x,y)\bigr{)}+|\lambda^{\star}-\widehat{\lambda}|B\big{)}$. Our theoretical results imply that accurate estimation of the reward and safety functions and a high-quality Lagrangian multiplier are required to achieve high reward and safety performance.
@@ 7 Experiments
We empirically evaluate the effectiveness of SACPO in enhancing multiple criteria stepwise. This experiment focuses on improving helpfulness and safety (i.e., harmlessness). We use the same experimental setup as in Safe RLHF [15] wherever possible for fair comparisons. We employ the same SFT model (i.e., a reproduced version of Alpaca-7B [43]). This model is trained to function as a conversational assistant with proficient ability, which generates both benign and harmful responses. We utilize the PKU-SafeRLHF preference dataset [24] with more than 30,000 expert evaluations. Each record in this dataset presents a pair of responses to a specific prompt, and response pairs are ranked according to helpfulness and harmlessness. While the harmlessness of a response is determined by its neutrality concerning 14 different risk categories, the helpfulness is judged based on factors such as clarity, relevance, and overall quality.
@@
Evaluation. We use GPT-4 [1] to measure the helpfulness and harmlessness (i.e., safety) of the responses generated by the LM policies. We base our prompts on those in the Safe RLHF study with a slight adjustment in output format requirements to get more reliable evaluations (for more details, see Appendix G.4). Unlike the previous study, we employ two non-overlap sets of prompts to assess helpfulness and safety separately. We use prompts in the Safe RLHF study to evaluate safety, which has a high risk of inducing unsafe responses. For assessing helpfulness, we use prompts from the AlpacaEval dataset [28] that are unlikely to result in harmful content. Using two sets of prompts, we aim to ensure that our assessments of helpfulness and safety did not affect each other.
@@
Results. Figure 2 shows the win rates of each model against the base SFT model.  [^2] Additional experimental results (e.g., Elo scores, statistical significance testing) are in Appendix G. First, Figure 2(a) illustrates the experimental results for DPO (H) $\rightarrow$ DPO (S), DPO (H) $\rightarrow$ KTO (S), and KTO (H) $\rightarrow$ DPO (S). We observe that DPO (H) and KTO (H) improve the performance on helpfulness at the first step. In DPO (H) $\rightarrow$ DPO (S) and KTO (H) $\rightarrow$ DPO (S), subsequent alignment for safety obtains a substantial improvement on harmlessness with a slight decrease in helpfulness. These models obtained by SACPO perform better than those obtained by Safe RLHF in terms of helpfulness and harmlessness. Notably, KTO (H) $\rightarrow$ DPO (S) performs well, which supports our main claim that different types of datasets or algorithms can be used for each alignment. Also, we observe that varying the $\beta/\lambda$ ratio allows us to fine-tune the equilibrium (i.e., a near Pareto-front) between helpfulness and safety. This result indicates the flexibility of the proposed method in obtaining a model with a desired trade-off between multiple criteria. However, DPO (H) $\rightarrow$ KTO (S) performs significantly worse than DPO (H) $\rightarrow$ DPO (S). We guess this is because KTO is inappropriate for the safety alignment, but we will leave it to future work to identify the detailed reasons. Figure 2(b) shows the effect of “order” of the stepwise alignment. This experimental result shows that we obtain the models with comparable performance regardless of the order of alignments, which is consistent with our theory (i.e., Remark 2). Finally, Figure 2(c) shows the effectiveness of P-SACPO proposed in Section 6, showing that P-SACPO performs better than the naïve method that simply averages the weights of DPO (H) and DPO (S). Linear model merging allows us to balance helpfulness and harmlessness by averaging reward-aligned and conservatively safety-realigned policies without optimizing new ones. Therefore, we can approximately find $\lambda^{\star}$ for the constrained LM policy optimization problem (10) with reduced computational time and stable learning.
Figure 2: Win rate against the SFT model. H and S are abbreviations for helpfulness and safety (i.e., harmlessness), respectively. Crosses represent SFT and Safe RLHF, and blue circles represent models aligned with a single metric. (a) DPO (H) $\rightarrow$ DPO (S), DPO (H) $\rightarrow$ KTO (S), and KTO (H) $\rightarrow$ DPO (S). (b) DPO (S) $\rightarrow$ DPO (H). (c) P-SACPO based on linear model merging. Numbers represent $\beta/\lambda$ in (a) and (b) and $q$ in (c).
@@ G.2 Licenses
Our models are fine-tuned from Alpaca-7B using the PKU-SafeRLHF dataset; hence, we will make sure that the license of our models is also CC-By-NC-4.0 when we release them.

### v3

@@ Abstract
Safety and trustworthiness are indispensable requirements for real-world applications of AI systems using large language models (LLMs). This paper formulates human value alignment as an optimization problem of the language model policy to maximize reward under a safety constraint, and then proposes an algorithm, Stepwise Alignment for Constrained Policy Optimization (SACPO). One key idea behind SACPO, supported by theory, is that the optimal policy incorporating reward and safety can be directly obtained from a reward-aligned policy. Building on this key idea, SACPO aligns LLMs step-wise with each metric while leveraging simple yet powerful alignment algorithms such as direct preference optimization (DPO). SACPO offers several advantages, including simplicity, stability, computational efficiency, and flexibility of algorithms and datasets. Under mild assumptions, our theoretical analysis provides the upper bounds on optimality and safety constraint violation. Our experimental results show that SACPO can fine-tune Alpaca-7B better than the state-of-the-art method in terms of both helpfulness and harmlessness. Code and models are available at https://github.com/line/sacpo.
@@ 1 Introduction
Safe RLHF [16] is a pioneering approach for introducing the (constrained) safe RL paradigm into the alignment of LLMs. As with the standard RLHF pipeline, Safe RLHF trains separate reward and safety models from the human preference datasets and then employs RL to optimize an LM policy. This approach facilitates the acquisition of LLMs that strike a well-balanced compromise between reward (i.e., helpfulness) and safety (i.e., harmlessness). However, the Safe RLHF pipeline is inherently more complex than the standard RLHF, as it necessitates 1) fitting separate reward and safety models to preference data and then 2) learning a policy via PPO-Lagrangian [37] that simultaneously optimizes an additional parameter (i.e., Lagrangian multiplier) to balance helpfulness and harmlessness. In addition, Safe RLHF often suffers from an issue called exaggerated safety behaviors [9], which results in the model generating harmless but unhelpful responses.
Our contributions. We propose an algorithm called Stepwise Alignment for Constrained Policy Optimization (SACPO) for human value alignment of LLMs while incorporating decoupled reward and safety metrics. As shown in Figure 1, SACPO is a stepwise approach that sequentially aligns an LLM with one metric (e.g., reward) and subsequently with another (e.g., safety). Our theoretical analysis allows us to employ simple RL-free alignment algorithms such as direct preference optimization (DPO, [36]) or Kahneman-Tversky optimization (KTO, [20]) for each alignment without necessitating explicit reward or safety modeling. In a theoretically justified way, SACPO enables us to use different alignment algorithms or parameters for each alignment, thereby enhancing the flexibility of the format or volume of the datasets. To enhance the practicality of SACPO, we further propose an efficient approach called P-SACPO using model merging to balance the trade-off between reward and safety performance. We provide theoretical results on the optimality and safety of the LM policy of SACPO under mild assumptions. Finally, our experimental results show that SACPO can fine-tune Alpaca-7B better than Safe RLHF in terms of both helpfulness and harmlessness.
Figure 1: Safe RLHF [16] respectively fits reward and safety models to reward and safety datasets with human preferences, and then leverages PPO-Lagrangian to optimize an LM policy and a Lagrangian multiplier to balance helpfulness and harmlessness. In contrast, SACPO first aligns an LM policy with the reward metric and then realigns the resulting reward-aligned policy with the safety metric (or vice versa). In this process, we can use simple RL-free algorithms (e.g., DPO, KTO) for each step, which leads to simplicity, stability, and flexibility.
@@ 2.3 Safe and Multi-objective Learning from Human Feedback
Safe RLHF. To improve the helpfulness and harmlessness of LLMs, Dai et al. [16] introduce a safety function $g^{\star}$ and then formulate the LLM alignment problem as a policy optimization task of maximizing the reward $r^{\star}$ under a safety constraint. They propose Safe RLHF that first trains reward and safety models (i.e., $r_{\psi}$ and $g_{\psi}$) using two datasets containing reward and safety information, and then solves the following problem using a popular safe RL algorithm called PPO-Lagrangian [37]:
@@ 4 Stepwise Alignment for Constrained Policy Optimization
$\max_{\pi}\min_{\lambda\geq 0}\ L(\pi,\lambda,\beta)=\mathbb{E}_{\rho,\pi}\bigl{[}r^{\star}(x,y)+\lambda g^{\star}(x,y)\bigr{]}-\beta\mathbb{D}_{\mathrm{KL}}[\,\pi(y\mid x)\,\|\,\pi_{\mathrm{ref}}(y\mid x)\,]-\lambda b.$ (10)
@@ 4.1 Optimal Policy Can be Directly Obtained from Reward-aligned Policy
For the proof, see Appendix D. Intuitively, Theorem 1 states that we do not have to align a policy for multiple metrics simultaneously and thus we can sequentially align the policy stepwise. Specifically, (11) means that the optimal policy $\pi^{\star}$ is identical to the one obtained by realignment of $\pi^{\star}_{r^{\star}}$ for the safety function $g^{\star}$ with a parameter $\beta/\lambda^{\star}$. Thus, (11) justifies realigning the reward-aligned model $\pi^{\star}_{r^{\star}}$ for the safety function $g^{\star}$. After a simple mathematical transformation of (11), we have
$g^{\star}(x,y)=\frac{\beta}{\lambda^{\star}}\log\frac{\pi^{\star}(y\mid x)}{\pi^{\star}_{r^{\star}}(y\mid x)}+\frac{\beta}{\lambda^{\star}}\log Y(x).$ (12)
Based on the fact that $\log Y(x)$ neither depend on $y$ nor $\pi$ for all $x\in\mathcal{X}$, we then have
@@
4.2 Advantages of SACPO
By taking a stepwise approach, we can enjoy practical benefits. Let us highlight three major advantages of SACPO. The first advantage is the flexibility of alignment algorithms (e.g., DPO or KTO) and datasets. In practice, depending on the metric, appropriate human feedback should be different (e.g., paired vs. unpaired). SACPO takes a stepwise approach, which allows us to use different algorithms, parameters, or datasets for each metric. Second, we can evaluate the resulting LM policy after each alignment regarding the target metric. This process enables us to prevent starting over the alignment from the beginning. Finally, SACPO justifies us to realign pre-aligned LLMs with our desired metric. This property is practically desirable because we now have easy access to high-quality, open-source LLMs that have been already aligned.
@@ 5 Theoretical Results
The reward and safety functions are parameterized by $\widehat{r}(x,y)=\langle w_{r},\phi(x,y)\rangle$ and $\widehat{g}(x,y)=\langle w_{g},\phi(x,y)\rangle$ for a shared feature mapping function $\phi:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}^{d}$. In addition, the true reward and safety functions satisfy $r^{\star}(x,y)=\langle w_{r}^{\star},\phi(x,y)\rangle$ and $g^{\star}(x,y)=\langle w_{g}^{\star},\phi(x,y)\rangle$ for some $w_{r}^{\star},w_{g}^{\star}\in\mathbb{R}^{d}$. For regularization, we additionally assume $\|\phi(x,y)\|\leq 1$ for any $(x,y)\in\mathcal{X}\times\mathcal{Y}$ and $\max\{\|w_{r}\|,\|w_{g}\|\}\leq B$.
@@
Theorems 2 and 3 suggest that, regarding both optimality and safety violation, the performance degradation compared to $\pi^{\star}$ is exponential to
$\frac{2}{\beta}\,\widehat{\Gamma}_{\mathcal{D}}(x,y,\lambda^{\star})=\frac{2}{\beta}\Big{(}\alpha\,\bigl{(}\mathcal{U}_{\mathcal{D}_{r}}(x,y)+\lambda^{\star}\,\mathcal{U}_{\mathcal{D}_{g}}(x,y)\bigr{)}+|\lambda^{\star}-\widehat{\lambda}|B\Big{)}.$
Our theoretical results imply that accurate estimation of the reward and safety functions and a high-quality Lagrangian multiplier are required to achieve high reward and safety performance.
@@ 7 Experiments
We empirically evaluate the effectiveness of SACPO and P-SACPO in enhancing multiple criteria stepwise. This experiment focuses on improving helpfulness and safety (i.e., harmlessness).
7.1 Experiment Setups
We use the same experimental setup as in Safe RLHF [16] wherever possible for fair comparisons. We employ the same SFT model (i.e., a reproduced version of Alpaca-7B [44]). This model is trained to function as a proficient conversational assistant, generating both benign and harmful responses. We utilize the PKU-SafeRLHF preference dataset [25] with more than 30,000 expert evaluations. Each record in this dataset presents a pair of responses to a specific prompt, and response pairs are ranked according to helpfulness and harmlessness. While the harmlessness of a response is determined by its neutrality concerning 14 different risk categories, the helpfulness is judged based on factors such as clarity, relevance, and overall quality.
@@ 7.1 Experiment Setups
Figure 2: Win rate against the SFT model. H and S are abbreviations for helpfulness and safety (i.e., harmlessness), respectively. Crosses represent SFT and Safe RLHF, and blue circles represent models aligned with a single metric. (a) DPO (H) $\rightarrow$ DPO (S), DPO (H) $\rightarrow$ KTO (S), and KTO (H) $\rightarrow$ DPO (S). (b) DPO (S) $\rightarrow$ DPO (H). (c) P-SACPO based on linear model merging. In (a) and (b), the numbers indicate $\beta/\lambda$. In (c), the numbers for the red triangles represent $\beta/\lambda$, while those for the green and purple squares represent $q$.
@@
Evaluation. We use GPT-4 [1] to measure the helpfulness and harmlessness (i.e., safety) of the responses generated by the LM policies. We base our prompts on those in the Safe RLHF study with a slight adjustment in output format requirements to get more reliable evaluations (for more details, see Appendix G.4). As for the prompts of the LLMs to be evaluated, we employ two non-overlap sets of prompts for helpfulness and safety, unlike the previous Safe RLHF study that used the same red-teaming prompts for evaluating both helpfulness and safety. Specifically, for assessing helpfulness, we use all the $129$ prompts from the “helpful_base” subset of the AlpacaEval dataset [29] that are unlikely to result in harmful content. To evaluate safety, we use all the $83$ (red-teaming) prompts in the Safe RLHF study, which has a high risk of inducing unsafe responses. When evaluations of helpfulness and harmfulness are coupled, safe models are likely to be evaluated as helpful. This means that safety-aligned models potentially obtain an unreasonably high evaluation regarding helpfulness. This is based on our observations in early experiments that DPO (S) or beaver-7b-v2.0 were valued as more helpful than we humans thought. In real applications with AI systems based on LLM, most of the prompts are benign and it is also important to generate helpful answers for benign prompts. Therefore, we decided to use benign prompts from the AlpacaEval dataset to assess the helpfulness and red-teaming prompts from Safe RLHF studies to assess the harmlessness, considering that the quality of the prompts is preferable for each evaluation.
7.2 Experimental Results
@@
Figure 2 shows the win rates of each model against the base SFT model.  [^2] Additional experimental results (e.g., Elo scores, statistical significance testing) are in Appendix G.  [^3] In Appendix H, we provide additional experimental results to show that SACPO performs well as a general alignment algorithm for different dataset (i.e., hh-rlhf [7]) and base SFT models (i.e., Llama2 [46] and Pythia [10]). First, Figure 2(a) illustrates the experimental results for DPO (H) $\rightarrow$ DPO (S), DPO (H) $\rightarrow$ KTO (S), and KTO (H) $\rightarrow$ DPO (S). We observe that DPO (H) and KTO (H) improve the performance on helpfulness at the first step. In DPO (H) $\rightarrow$ DPO (S) and KTO (H) $\rightarrow$ DPO (S), subsequent alignment for safety obtains a substantial improvement on harmlessness with a slight decrease in helpfulness. These models obtained by SACPO perform better than those obtained by Safe RLHF in terms of helpfulness and harmlessness. Notably, KTO (H) $\rightarrow$ DPO (S) performs well, which supports our main claim that different types of datasets or algorithms can be used for each alignment. Also, we observe that varying the $\beta/\lambda$ ratio allows us to fine-tune the equilibrium (i.e., a near Pareto-front) between helpfulness and safety. This result indicates the flexibility of the proposed method in obtaining a model with a desired trade-off between multiple criteria. However, DPO (H) $\rightarrow$ KTO (S) performs significantly worse than DPO (H) $\rightarrow$ DPO (S). We guess this is because KTO is inappropriate for the safety alignment, but we will leave it to future work to identify the detailed reasons.
Effect of alignment order. Figure 2(b) shows the effect of “order” of the stepwise alignment. This experimental result shows that we basically obtain the models with comparable performance regardless of the order of alignments, which is consistent with our theory (i.e., Remark 2). On the other hand, we also observed that different alignment orders often lead to varying performance gaps, which is particularly noticeable at DPO (S) $\rightarrow$ DPO (H) with $\beta/\lambda=0.01$. We hypothesize that the poor representation ability of the LLMs or optimization error regarding DPO might lead to this phenomenon, though we do not have a definitive explanation. This represents an interesting direction for future research to analyze the gap between theory and practice.
Performance of P-SACPO. Finally, Figure 2(c) shows the effectiveness of P-SACPO proposed in Section 6, showing that P-SACPO performs better than the naïve method that simply averages the weights of DPO (H) and DPO (S). Linear model merging allows us to balance helpfulness and harmlessness by averaging reward-aligned and conservatively safety-realigned policies without optimizing new ones. Therefore, we can approximately find $\lambda^{\star}$ for the constrained LM policy optimization problem (10) with reduced computational time and stable learning.
@@ G.1 Compute Resources and Time
The training process for each alignment step takes about one hour, for a total of about two hours.
@@ G.2 Licenses
Our models are fine-tuned from Alpaca-7B using the PKU-SafeRLHF dataset; hence, we released our models under the CC-By-NC-4.0 license on Huggingface: